# Natural, Artificial, and Human Intelligences

### v1

@@ 8 Our Proposal
Chatbots have fundamentally challenged our notion of intelligence. One reason is the role of (especially written) language in human intelligence and civilization. Language has been a key facilitator of progress in our species, by allowing inventions which persist beyond the individual, collaboration/ role differentiation, and communities of knowledge (Fernbach and Sloman,, 2018). Historically, human linguistic ability has been pillar in our self-concept of intelligence and assumption of superiority — so, with the emergence of chatbots, an observer might say that we have to either declare chatbots intelligent or argue that being a wordsmith is no longer enough and change the rules for determining intelligence that we have insisted on for centuries.
We think there are some aspects of the intelligence debate which can be resolved fairly simply. To summarise, first, intelligence is more than just skill, however exceptional, whether this skill concerns arithmetic ability (we do not call a calculator intelligent), agility in some sport, or language (Chollet,, 2019). A conundrum is that, for example, a human who is exceptionally good at arithmetic might be called intelligent, but not a machine designed for this purpose. The resolution concerns generality of scope: intelligence implies a degree of generality, adaptability, and transferability. Regarding chatbots, if they were exceptionally good just in language, we would be (more) reluctant to decide that they are intelligent; the evidence suggests that their capability exceeds language (e.g., (xAI Team,, 2025)). Second, there is a criticism that we should not talk about intelligence before understanding what intelligence is in the first place. It would help to have such an understanding. However, there is a long list of concepts we have studied successfully prior to understanding what they are, such as life, love, space, gravitation etc. If we agree that certain examples of behaviour reflect intelligence, we can use these examples as a guide of when we have intelligence. Third, speaking of such outcomes, the ultimate outcomes of human intelligence are our scientific, cultural, and artistic achievements; so we can use these as a guide regarding whether chatbots are intelligent.
As Turing, (1950) noted, there may be a problem that, in assessing AI against such outcomes, we are comparing average AI against the very best of humans. We think the issue is what human intelligence is capable of, rather than the intelligence of the average human. Indeed, in studies of animal intelligence a similar approach has been adopted (e.g., (Range et al.,, 2011), (Güntürkün et al.,, 2024)). It is also true that a huge enabling factor for even the most celebrated scientists, artists etc. is the vast community of knowledge we have, largely thanks to written language — as Newton famously said, ”if I have seen further, it is by standing on the shoulders of giants.” However, we can recognise intelligence (for both humans and otherwise) both at the individual level and species level. Most organisms benefit from a degree of evolutionary intelligence, which often leads to amazing competencies – humans benefit from evolutionary intelligence too, however, with the advent of written language, the contribution from communities of knowledge has been far important.
There are a few more issues which we think are straightforward. Should it matter how an intelligent behaviour is produced, before we characterise an animal/ system etc. as intelligent? For example, does it matter whether similar priors were employed in solving a problem (Chollet,, 2019)? We think not, because the means of achieving an intelligent behaviour would depend on many factors, including hardwired vs. learnt structures, embodiment constraints, collaboration opportunity with other agents etc. Also, relatedly, does it matter whether some behaviour is accomplished more efficiently by one agent vs. less so by another? For similar reasons, we think not.
With these thoughts in mind, our proposal is as follows. We suggest that intelligence entails invention, which presupposes the ability to learn new skills, imagination/ creativity, hypothesis testing including trial and error (the ability to turn error and novelty into opportunity), and generalisation. Consider penicillin: the basic discovery was a matter of trial and error, but then Fleming’s genius as an inventor came from imagining how useful that error could be: this is an example of invention combining trial and error, hypothesis making, and generalisation (Fleming,, 1929). Intelligence also requires capacity for complex inference, knowledge, and the ability to create general thoughts. That is, intelligent agents must have some mechanism for organising elementary thoughts and pieces of information into a complex narrative. The catalytic role of language is yet again apparent, since it provides a fairly precise way to organise and work on large amounts of knowledge. Next, if we consider some of the most impressive accomplishments of human intelligence, such as the creation of scientific theories, would the above suffice? To develop such theories, we think that an agent needs to have awareness of its own understanding and so how this understanding is limited; for this to be possible, a prerequisite is a sense that the agent is distinguished from its environment. That is, together with the above, we think that the process of doing (notably) science is self-referential: it relates to what we do not understand and what we would like to understand/ how we can go beyond our current knowledge. Therefore, we suggest that an intelligent agent has to be self-aware and understand the distinction between itself and its environment. Finally, we think that intelligence requires embodiment and knowledge of its embodiment. For most agents, embodiment is what drives behaviour, because embodiment creates needs for resources. Also, embodiment constrains the means of acquiring knowledge about the environment and provides ground truths.
[…]
What about chatbots? They share this powerful tool with us, but they seem to lack some of the other characteristics of intelligence. They lack knowledge of their embodiment and so they have no needs and no curiosity — this limits what they could be doing. They are not self-aware (that we know of) and so this limits their capacity to reflect on what they know and what they could be trying to discover. However, they seem to be capable of inference, as is evidenced by their ability to follow through complex arguments. Regarding whether they are capable of invention, when chatbots produce something new, we tend to say they are hallucinating. Is this unfair dismissal of what might be capacity for creativity? The boundaries between imagination, creativity, and just getting things wrong are difficult to pin down. Nevertheless, for facts which can be easily verified and without insight of diverging from some established belief, we think it would be hard to justify a statement as creative vs. just wrong and chatbot hallucinations seem to be instances of the latter. We apply a similar criterion to human thought, for example, false memories in legal proceedings are not considered instances of creativity (e.g., Howe, (2013)); however, using the term ’hallucinations’ for chatbots is alarmist and obscures the fact that there similar human behaviours.
This does not mean that our jobs are safe from chatbots: chatbots could replace humans in any task which is reduced to wordsmithing and the simple but somewhat embarrassing truth is that unique human competencies invariably correspond to the animal stuff — tasks in which we excel because of our particular embodiment and for which a human empathic connection is useful. This is not necessarily a bad thing: perhaps if we, as a species, revised our perception of how unique our intelligence is, we might approach each other, our environment, and other species with a bit more humility. Equally, the idea that just training bigger and better chatbots will eventually lead us to true Artificial General Intelligence is probably wrong: insofar as rapid symbol-manipulation and enormous memory are part of intelligence, chatbots already surpass human capabilities, but also demonstrate that there is more to intelligence than the heavily-emphasized ability to write intelligent words.

### v2

@@ 8 Concluding Comments
Chatbots have fundamentally challenged our notion of intelligence. One reason is the role of (especially written) language in human intelligence and civilization. Language has been a key facilitator of progress in our species, by allowing inventions which persist beyond the individual, collaboration/ role differentiation, and communities of knowledge (fernbach2018knowledge). Historically, human linguistic ability has been pillar in our self-concept of intelligence and assumption of superiority — so, with the emergence of chatbots, an observer might say that we have to either declare chatbots intelligent or argue that being a wordsmith is no longer enough, and change the rules for determining intelligence that we have insisted on for centuries.
Given the prowess of chatbots on IQ tests and their many offshoots, it is tempting to look for definitions of intelligence that emphasize other aspects of human achievement. Even chatbots designed for coding often lack the ability to observe what’s going on in different parts of a computer screen, and they can propose (and sometimes execute) syntactically-correct changes to code that have alarmingly bad consequences (kim2025vibe). It is thus tempting to argue that AI still lack ‘general human-level intelligence’, on the grounds that some of the most advanced systems available clearly lack some of the basic physical and social intelligences observed in many animals.
However, while such arguments have been used to explain the historic success of humans in terms of ‘superior intelligence’, they are likely to become obsolete due to progress in robotics and cooperative agents. The potential for incorporating large language models in embodied agents is demonstrated by authors including zhang2023building, and robots that dance and perform typical human physical activities are already being marketed for consumer use. It is easy to anticipate machines that will win at tennis before long, in the same way that, by the early 1990s, it had become easy to anticipate that sooner or later machines would win at chess (kasparov2017deep). The prototypical evolutionary example of human prowess in embodied and social intelligence might be reflected in the collaboration in a hunting or war party. However, it is hard to imagine that even the best human commandos would defeat a fleet of military drone hunters within a few years. We may object to armed drones on moral grounds, but semi-autonomous armed drones and components are already commercially available (summerfield2024strange).
On the initial question, “Are humans uniquely intelligent?”, chatbots demonstrate that humans are no longer unique at demonstrating the symbol-processing intelligences that have been prized and measured since Victorian times. While this conclusion does not contradict any established scientific test, it is disruptive to the scientific establishment.
It is nonetheless obvious that chatbots alone do not demonstrate Artificial General Intelligence (AGI), in the sense of performing all cognitive tasks at least as well as a human — in 2025, glaring common sense blunders are still far too prevalent. For some years, training bigger and better chatbots led to such improvements that the approximation to human conversational intelligence is often surprisingly good (summerfield2024strange), but this trend has already been superseded: advances in 2025 have highlighted reasoning models, and frontier tasks now include semantically-accurate video generation.
This widely accepted definition of AGI is explicitly anthropocentric. Should humans be accepted as a general standard of intelligence for all cognitive tasks? Even by 1950, Turing observed that humans no longer had any hope of performing arithmetic as fast as machines: so if arithmetic is one of the cognitive tasks, humans cannot hope to surpass the ‘general intelligence’ of machines. Key advances in the scientific understanding of space and time have come from realising that man is not the measure of all things, and human scales are small compared to the universe. Truly general intelligence need not be limited by the intelligence of a human; we hope that allowing ourselves to distinguish these will eventually help us to understand intelligence and humans better.
It increasingly appears that chatbots could replace humans in any task which is reduced to wordsmithing and the simple but somewhat embarrassing truth is that unique human competencies correspond to the animal stuff — tasks in which we excel because of our particular embodiment, and for which a human empathic connection is useful. However, given the rapid development of multimodal AI models and semi-autonomous robots, this assessment may become dated within a few years. Of course, replacing tasks does not always equate to replacing jobs: in some professions, including medicine and social work, reducing the burden of paperwork is a goal in itself (johnson2021electronic). This is not necessarily a bad readjustment: perhaps if we, as a species, revised our perception of how unique our intelligence is, we might approach each other, our environment, and other species with a bit more humility.